# Closed-loop Long-horizon Robotic Planning via Equilibrium Sequence Modeling

### v4

@@ Abstract
In the endeavor to make autonomous robots take actions, task planning is a major challenge that requires translating high-level task descriptions into long-horizon action sequences. Despite recent advances in language model agents, they remain prone to planning errors and limited in their ability to plan ahead. To address these limitations in robotic planning, we advocate a self-refining scheme that iteratively refines a draft plan until an equilibrium is reached. Remarkably, this process can be optimized end-to-end from an analytical perspective without the need to curate additional verifiers or reward models, allowing us to train self-refining planners in a simple supervised learning fashion. Meanwhile, a nested equilibrium sequence modeling procedure is devised for efficient closed-loop planning that incorporates useful feedback from the environment (or an internal world model). Our method is evaluated on the VirtualHome-Env benchmark, showing advanced performance with better scaling for inference computation. Code is available at https://github.com/Singularity0104/equilibrium-planner.
@@ 1 Introduction
Recent advances in large language models (LLMs) have spurred tremendous progress in robotic planning (Huang et al., 2022; Li et al., 2022; Singh et al., 2023; Driess et al., 2023; Ahn et al., 2023; Huang et al., 2023; Zhao et al., 2023; Hu et al., 2024). Based on their extensive world knowledge, LLM agents seem close to autonomously performing robotic tasks, such as in household scenarios. However, growing evidence shows that existing LLM agents struggle with task planning (Kaelbling & Lozano-Pérez, 2011) that decomposes a high-level task into mid-level actions. While this problem requires long-horizon planning as well as consideration of environmental feedback, LLMs are often limited by: (1) unidirectional dependency: due to autoregressive generation, previous tokens cannot attend to future tokens, resulting in limited ability to plan ahead (Wu et al., 2024a); (2) lack of error correction for existing outputs, unless with a heavy system 2; (3) fixed forward process hindering the allocation of more inference computation to further improve planning performance. These inherent limitations of LLMs lead to inefficiency in the closed-loop long-horizon robotic planning.
To address above challenges of LLM planners in closed-loop long-horizon planning, we advocate the approach of self-refinement (Welleck et al., 2023; Shinn et al., 2023; Kim et al., 2023; Madaan et al., 2023) that iteratively improves a previously generated plan. The reasons behind are threefold: (1) bidirectional dependency: since the output is conditioned on a previous draft plan, it can attend to all tokens in the plan (from an old version), thus improving its ability to plan ahead; (2) internal error correction which allows implicit self-correction in a forward pass without an explicit, heavy system 2; (3) dynamic computation allocation by iterating through a self-refinement process until convergence. However, such a self-refining strategy imposes significant training difficulties because it requires backpropagation through infinite self-refining steps (Werbos, 1990). This may be seen as an extreme case that reflects some of the general challenges in teaching LLMs to plan and reason. Existing solutions include curating process supervision (Uesato et al., 2022; Lightman et al., 2024) or applying reinforcement learning (Zelikman et al., 2024; OpenAI, 2024; Kumar et al., 2024), but they are considerably more complex than supervised training and difficult to implement.
This work proposes an simple learning framework for planning via self-refinement. Specifically, we formulate the self-refining process as a fixed-point problem that recursively refines the plan until the equilibrium point, as illustrated in Fig. 1. While the forward process of this fixed-point problem could be solved efficiently using root-finding methods, more interestingly, its backpropagation can be skipped since its gradient is explicated by the implicit function theorem (Krantz & Parks, 2002) as in deep equilibrium models (Bai et al., 2019; Geng & Kolter, 2023). It is noted that the derived gradient term may be further simplified through a Jacobian-free approximation (Fung et al., 2022) to facilitate training. These analytical techniques allow end-to-end supervised training of the LLM planner to accomplish self-refinement without the need for additional verifiers or reward models in reinforcement learning-based counterparts, greatly enhancing simplicity and practicality. And after training, our equilibrium model-based planner is capable of dynamically allocating more inference computation in the equilibrium solving process to achieve better planning performance.
Another important cue for self-refinement in robotic tasks is closed-loop feedback from the environment. To efficiently incorporate environmental feedback, we devise a nested equilibrium sequence modeling procedure consisting of inner and outer loops, where the inner loop iteratively refines a plan using previous feedback, while the outer loop updates the feedback by interacting with the environment. This enables closed-loop planning from even a few environmental interactions. Moreover, the nested equilibrium solving process is accelerated by reusing the previously derived equilibrium. We further implement the above design within an LLM agent framework, seamlessly integrating the equilibrium model-based planner, an experience memory buffer containing past plans and feedback, and a world model to estimate feedback in the absence of environmental interactions, thus enabling it to operate effectively in robotic planning. The contributions of our work are summarize as follows:
  • We present equilibrium sequence modeling, a simple training approach for self-refining LLMs based on equilibrium models, allowing for end-to-end learning in a supervised manner.
  • Our method is evaluated on the VirtualHome-Env benchmark (Puig et al., 2018; Liao et al., 2019), demonstrating its advantageous performance with better scaling w.r.t. inference computation.
@@
Figure 1: Illustration of the equilibrium point in planning. We view planning as a self-refinement process in which the ideal plan emerges as an equilibrium point, remaining unchanged by any refinement attempts even with newer information (e.g. feedback from the environment or a world model). This enables us to tackle robotic planning from an optimization perspective around its equilibrium.
@@ 2 Related Work
LLMs for Planning. Scaling up inference computation to improve LLMs’ performance on planning and reasoning tasks has received increasing attention (Brown et al., 2024; Snell et al., 2024; Wu et al., 2024b; OpenAI, 2024). Existing techniques involving chain-of-thought (Wei et al., 2022; Zelikman et al., 2022; 2024), repeated sampling (Wang et al., 2023; Brown et al., 2024) and tree search (Yao et al., 2023a; Zhao et al., 2023) showed preliminary results with handcraft system 2. Alternatively, a method called self-refinement (Welleck et al., 2023; Shinn et al., 2023; Kim et al., 2023; Madaan et al., 2023) suggests recursively refining the existing LLM output in an autonomous manner, but it relies heavily on prompting or intricate training procedures. To fully exploit its potential, we propose an end-to-end optimization method for self-refinement through deep equilibrium models.
Deep Equilibrium Models (Bai et al., 2019) are infinite-depth neural networks specified by fixed-point problems $x^{*}=f_{\theta}(x^{*})$, where $f_{\theta}$ is an equilibrium layer. While their inference can take infinite steps by the fixed-point iteration, their gradients are estimated using implicit differentiation (Krantz & Parks, 2002) without backpropagating through all layers, thus enabling memory-efficient training. They have been extensively applied to tasks such as visual understanding (Bai et al., 2020; 2022) and image generation (Pokle et al., 2022; Geng et al., 2023; Bai & Melas-Kyriazi, 2024). In this paper, we apply the fixed-point formulation of deep equilibrium models to the self-refinement process in LLM planners, allowing for simple supervised training of LLMs to refine themselves. More detailed introduction to deep equilibrium models is presented in Appendix A.
@@ 3 Method
We study the problem of robot task planning which aims to decompose a high-level task description into long-horizon mid-level action sequences (Kaelbling & Lozano-Pérez, 2011). In the following, we first discuss the limitations of LLM planners in self-refinement (Section 3.1) and address them with a novel equilibrium sequence modeling scheme (Section 3.2), which is compatible with various feedback from the designs in Section 3.3. Lastly, practical implementations are given in Section 3.4.
@@ 3.1 Preliminaries on Self-Refinement
The prevailing LLMs are intrinsically limited in planning, as their unidirectional dependency results in limited capability to plan ahead (Wu et al., 2024a), and the lack of error correction hinders closed-loop planning. These reasons call for alternative mechanisms to address robot task planning.
Recently, Welleck et al. (2023); Shinn et al. (2023); Kim et al. (2023); Madaan et al. (2023) proposed self-refinement, which uses an LLM $f_{\theta}$ to iteratively refine the previous LLM output. This strategy naturally addresses the above limitations, since it introduces bidirectional token dependency and a dynamic error correction mechanism. Formally, let $x_{t}$ denote a draft plan and $c_{t}$ denote context (e.g. environmental feedback), then planning may be viewed as a self-refinement process as follows:
@@
However, self-refinement via prompting (Shinn et al., 2023; Kim et al., 2023; Madaan et al., 2023) has been found to be very limited by Huang et al. (2024). Alternative training-based methods require careful curation of training sequences (Welleck et al., 2023; Havrilla et al., 2024) or reinforcement learning (Qu et al., 2024; Kumar et al., 2024) and are therefore difficult to train. Overall, they remain deficient for robotic planning compared to system 2-based alternatives, as shown in Hu et al. (2024).
@@ 3.2 Self-Refinement as An Equilibrium Model
Let us first consider a simplified scenario of self-refinement without environmental feedback, namely that the context $c_{t}$ is fixed, e.g. to a predefined system message $c$. Then, the self-refinement process in Eq. 1 reduces to a fixed-point problem concerning only the plan $x_{t}$. Denote the initial plan by $x_{0}=\varnothing$ and the equilibrium plan, i.e. the endpoint, by $x^{*}$, then its trajectory can be expressed as:
@@
Instead, we approach it directly from an analytical perspective. Assuming access only to an outcome supervision $L(\cdot,y)$ on the plan, e.g. its distance to the ground truth plan $y$, then self-refinement is formulated as an optimization problem minimizing the loss function of the equilibrium plan:
@@
Theorem 1
(Implicit Function Theorem (Bai et al., 2019; Krantz & Parks, 2002)) Assuming that $\big{(}I-\frac{\partial f_{\theta}}{\partial x^{*}}\big{)}$ is invertible, then the loss gradient of Eq. 3 w.r.t. $\theta$ is given by:
@@
Its proof is given in Section A.4. It is noteworthy that the inverse Jacobian term $A=(I-\frac{\partial f_{\theta}}{\partial x^{*}})^{-1}$ within the above gradient is difficult to compute exactly, for which existing work often approximates through the damped fixed-point unrolling or the Neumann series (Geng et al., 2021b). For computa-tional efficiency, we drop the inverse Jacobian term using $A\approx I$ as in Fung et al. (2022); Geng et al. (2021a); Choe et al. (2023), the latter having been validated on Transformer-based LLMs:
@@
Figure 2: Illustration of equilibrium sequence modeling. Prior to training, our model first undergoes iterative inference to reach an equilibrium plan. It is then pushed away from the equilibrium towards ground truth by a supervised training loss. This process enables self-refinement of the model.
@@
This new formula represents a new equilibrium sequence modeling scheme that can be implemented in two stages as in Fig. 2. In the first stage, we iterate over the fixed-point problem within Eq. 3 to solve the equilibrium plan $x^{*}$. Then, it is paired with the ground truth plan $y$ as a training sequence, which is used as in the standard supervised finetuning pipeline to teach the LLM to self-refine.
It features two intuitive advantages: (1) instead of directly regressing the ground truth, it only adjusts the equilibrium point, which reduces overfitting compared to the vanilla supervised finetuning; (2) it teaches self-refinement by a simple supervised loss, without requiring additional value functions or reward models (Welleck et al., 2023; Havrilla et al., 2024; Qu et al., 2024; Kumar et al., 2024).
@@ 3.3 Equilibrium Models with Feedback
To take into account environmental feedback, we consider an adaptive context $c_{t}$ that is influenced by the plan $x_{t}$ rather than fixed. Then, the previously concerned equilibrium solving process of Eq. 2 should be revised as an iterative process that couples plan and feedback, starting from $x_{0}=c_{0}=\varnothing$:
@@
After the modification, the existing derivations only hold when we neglect the derivatives related to $c^{*}$. Fortunately, this is a natural choice due to the non-differentiability of most feedbacks. Therefore, the equilibrium planner can be trained in a similar supervised way as in Eqs. 6 and 2, and after training it would be able to self-refine based on the latest feedback just by forward passes. However, iteratively interacting with the environment to obtain feedback is costly and may not be recoverable. In response, we devise a nested equilibrium solving scheme for more efficient closed-loop planning.
@@
planner $f_{\theta}$, environment or world model Env, number of iterations $N$.
Initialize a start point $x_{0}$ and feedback $c_{0}$.
for $i\leftarrow 0$ to $N$ or converged do
     Solve inner equilibrium loop to obtain $x_{t}^{*}$. $\triangleright$ Eq. 8
     Update next plan $x_{t+1}$ and feedback $c_{t+1}$ with Env. $\triangleright$ Eqs. 9 and 10
generated plan $x^{*}$.
@@
Thanks to this inner-loop introspection mechanism, our equilibrium model can be more efficient in closed-loop planning, achieving superior performance with fewer environmental interactions.
@@
which corresponds to the starting point $x_{t+1}^{0}$ of the inner loop. Similarly, history feedback could be reused across different inner loops. This is achieved by initializing the context of the next inner loop by concatenating the previous feedback with the latest feedback (paired with its original plan):
@@
where $\parallel$ denotes concatenation. The whole nested inference procedure is described in Algorithm 1.

### v5

@@ Abstract
In the endeavor to make autonomous robots take actions, task planning is a major challenge that requires translating high-level task descriptions to long-horizon action sequences. Despite recent advances in language model agents, they remain prone to planning errors and limited in their ability to plan ahead. To address these limitations in robotic planning, we advocate a self-refining scheme that iteratively refines a draft plan until an equilibrium is reached. Remarkably, this process can be optimized end-to-end from an analytical perspective without the need to curate additional verifiers or reward models, allowing us to train self-refining planners in a simple supervised learning fashion. Meanwhile, a nested equilibrium sequence modeling procedure is devised for efficient closed-loop planning that incorporates useful feedback from the environment (or an internal world model). Our method is evaluated on the VirtualHome-Env benchmark, showing advanced performance with improved scaling w.r.t. inference-time computation. Code is available at https://github.com/Singularity0104/equilibrium-planner.
Machine Learning, ICML
@@ 1 Introduction
Recent advances in large language models (LLMs) have spurred tremendous progress in robotic planning (Huang et al., 2022; Li et al., 2022; Singh et al., 2023; Driess et al., 2023; Ahn et al., 2023; Huang et al., 2023; Zhao et al., 2023; Hu et al., 2024). Based on their extensive world knowledge, LLM agents seem close to autonomously performing robotic tasks, such as in household scenarios. However, growing evidence shows that existing LLM agents struggle with task planning (Kaelbling & Lozano-Pérez, 2011) that decomposes a high-level task into mid-level actions. While this problem requires long-horizon planning as well as consideration of environmental feedback, LLMs are often limited by: (1) unidirectional dependency: due to autoregressive generation, previous tokens cannot attend to future tokens, resulting in limited ability to plan ahead (Wu et al., 2024); (2) lack of error correction for existing outputs, unless with a heavy system 2; (3) fixed forward process hindering the allocation of more inference computation to further improve planning performance. These inherent limitations of LLMs inhibit closed-loop long-horizon robotic planning.
@@
Figure 1: Illustration of the equilibrium point in planning. We view planning as a self-refinement process in which the ideal plan emerges as an equilibrium point, remaining unchanged by any refinement attempts even with newer information (e.g. feedback from the environment or a world model). This enables us to tackle robotic planning from an optimization perspective around its equilibrium, bypassing the need for sophisticated reinforcement learning.
To address the above challenges of LLM planners in closed-loop long-horizon planning, we advocate the approach of self-refinement (Welleck et al., 2023; Shinn et al., 2023; Kim et al., 2023b; Madaan et al., 2023) that iteratively improves a previously generated plan. The reasons behind are threefold: (1) bidirectional dependency: since the output is conditioned on a previous draft plan, it can attend to all tokens in the plan (from an old version), thus improving its ability to plan ahead; (2) internal error correction which allows implicit self-correction in a forward pass without an explicit, heavy system 2; (3) dynamic computation allocation by iterating through a self-refinement process until convergence. However, such a self-refining strategy imposes significant training difficulties because it requires backpropagation through infinite self-refining steps (Werbos, 1990). This may be seen as an extreme case that reflects some of the general challenges in teaching LLMs to plan and reason. Existing solutions include curating process supervision (Uesato et al., 2022; Lightman et al., 2024) or applying reinforcement learning (Zelikman et al., 2024; Jaech et al., 2024; Kumar et al., 2025; Guo et al., 2025), but they are considerably more complex than supervised training and remain unproven in the absence of efficient verifiers.
This work proposes a simple supervised learning framework for planning via self-refinement. Specifically, we formulate the self-refining process as a fixed-point problem that recursively refines the plan until the equilibrium point, as illustrated in Figure 1. While the forward process of this fixed-point problem could be solved efficiently using root-finding methods, more interestingly, its backpropagation can be skipped since its gradient is explicated by the implicit function theorem (Krantz & Parks, 2002) as in deep equilibrium models (Bai et al., 2019; Geng & Kolter, 2023). It is noted that the derived gradient term may be further simplified through a Jacobian-free approximation (Fung et al., 2022) to facilitate training. These analytical techniques allow end-to-end supervised training of the LLM planner to accomplish self-refinement without the need for additional verifiers or reward models in reinforcement learning-based counterparts, greatly enhancing simplicity and practicality. And after training, our equilibrium model-based planner dynamically allocates more inference computation based on the number of iterations needed to solve the equilibrium, thereby achieving better planning performance.
Another important cue for self-refinement in robotic tasks is closed-loop feedback from the environment. To efficiently incorporate environmental feedback, we devise a nested equilibrium sequence modeling procedure consisting of inner and outer loops, where the inner loop iteratively refines a plan using previous feedback, while the outer loop updates the feedback by interacting with the environment. This enables closed-loop planning from even a few environmental interactions. Moreover, the nested equilibrium solving process is accelerated by reusing the previously derived equilibrium. We further implement the above design within an LLM agent framework, seamlessly integrating the equilibrium model-based planner, an experience memory buffer containing past plans and feedback, and a world model to estimate feedback in the absence of environmental interactions, thus allowing the planning system to operate effectively in closed-loop long-horizon robot task planning scenarios. The core contributions of our work are as follows:
  • We present equilibrium sequence modeling, a simple training approach for self-refining LLMs based on equilibrium models, allowing for end-to-end supervised learning without additional verifiers or reward models.
  • Our method is evaluated on the VirtualHome-Env benchmark (Puig et al., 2018; Liao et al., 2019), demonstrating its advantageous performance with better scaling w.r.t. inference computation than tree-based alternatives.
@@ 2 Related Work
LLMs for Planning. Scaling up inference computation to improve LLMs’ performance on planning and reasoning tasks has received increasing attention (Brown et al., 2024; Snell et al., 2025; Wu et al., 2025; Jaech et al., 2024; Guo et al., 2025). Precedent techniques involving chain-of-thought (Wei et al., 2022; Zelikman et al., 2022, 2024), repeated sampling (Wang et al., 2023; Brown et al., 2024) and tree search (Yao et al., 2023a; Zhao et al., 2023) showed preliminary results with handcraft system 2. Alternatively, a method called self-refinement (Welleck et al., 2023; Shinn et al., 2023; Kim et al., 2023b; Madaan et al., 2023) suggests recursively refining the existing LLM output in an autonomous manner, but it relies heavily on prompting or sophisticated reinforcement learning. To fully exploit its potential, we propose an end-to-end optimization method for self-refinement via deep equilibrium models.
Deep Equilibrium Models (Bai et al., 2019) are infinite-depth neural networks specified by fixed-point problems $x^{*}=f_{\theta}(x^{*})$, where $f_{\theta}$ is an equilibrium layer. While their inference can take infinite steps by the fixed-point iteration, their gradients are estimated using implicit differentiation (Krantz & Parks, 2002) without backpropagating through all layers, thus enabling memory-efficient training. They have been extensively applied to tasks such as visual understanding (Bai et al., 2020, 2022) and image generation (Pokle et al., 2022; Geng et al., 2023; Bai & Melas-Kyriazi, 2024). In this paper, we apply the fixed-point formulation of deep equilibrium models to the self-refinement process in LLM planners, allowing for simple supervised training to refine themselves. More detailed introduction to deep equilibrium models is presented in Appendix A.
@@ 3 Method
We study the problem of robot task planning which aims to decompose a high-level task description into long-horizon mid-level action sequences (Kaelbling & Lozano-Pérez, 2011). In the following, we first discuss the limitations of LLM planners in self-refinement (Section 3.1) and address them with a novel equilibrium sequence modeling scheme (Section 3.2). This framework is adapted to closed-loop feedback with efficient designs in Section 3.3. Finally, practical implementations are presented in Section 3.4.
@@ 3.1 Preliminaries on Self-Refinement
The prevailing LLMs are intrinsically limited in planning, as their unidirectional dependency results in limited capability to plan ahead (Wu et al., 2024), and the lack of error correction hinders closed-loop planning. These reasons call for alternative mechanisms to address robot task planning.
Recently, Welleck et al. (2023); Shinn et al. (2023); Kim et al. (2023b); Madaan et al. (2023) proposed self-refinement, which uses an LLM $f_{\theta}$ to iteratively refine the previous LLM output. This strategy naturally addresses the above limitations, since it introduces bidirectional token dependency and a dynamic error correction mechanism. Formally, let $x_{t}$ denote a draft plan and $c_{t}$ denote context (e.g. environmental feedback), then planning may be viewed as a self-refinement process as follows:
@@
However, self-refinement via prompting (Shinn et al., 2023; Kim et al., 2023b; Madaan et al., 2023) has been found to be very limited by Huang et al. (2024). Alternative training-based methods require careful curation of training sequences (Welleck et al., 2023; Havrilla et al., 2024) or reinforcement learning (Qu et al., 2024; Kumar et al., 2025) and are therefore difficult to implement, even for domains with efficient verifiers such as coding and math. Overall, they remain deficient for robotic planning compared to system 2-based alternatives, as shown in Hu et al. (2024).
@@ 3.2 Self-Refinement as An Equilibrium Model
Let us first consider a simplified scenario of self-refinement without environmental feedback, namely that the context $c_{t}$ is fixed, e.g. to a predefined system message $c$. Then, the self-refinement process in Equation 1 reduces to a fixed-point problem concerning only the plan $x_{t}$. Denote the initial plan by $x_{0}=\varnothing$ and the equilibrium plan, i.e. the endpoint, by $x^{*}$, then its trajectory can be expressed as:
@@
Instead, we approach it directly from an analytical perspective. Assuming access only to outcome supervision $L(\cdot,y)$ on the plan, e.g. its distance to the ground truth plan $y$, then self-refinement is formulated as an optimization problem minimizing the loss function of the equilibrium plan:
@@
Theorem 3.1.
(Implicit Function Theorem (Bai et al., 2019; Krantz & Parks, 2002)) Assuming that $\big{(}I-\frac{\partial f_{\theta}}{\partial x^{*}}\big{)}$ is invertible, then the loss gradient of Equation 3 w.r.t. $\theta$ is given by:
@@
Its proof is given in Section A.4. It is noteworthy that the inverse Jacobian term $A=(I-\frac{\partial f_{\theta}}{\partial x^{*}})^{-1}$ within the above gradient is difficult to compute exactly, for which existing work often approximates through the damped fixed-point unrolling or the Neumann series (Geng et al., 2021b). For computational efficiency, we drop the inverse Jacobian term using $A\approx I$ as in Fung et al. (2022); Geng et al. (2021a); Choe et al. (2023), the latter work having been validated on Transformer-based LLMs:
@@
For a comprehensive introduction to equilibrium models and this approximation, please refer to Appendix A.
@@
Figure 2: Illustration of equilibrium sequence modeling with two alternating steps: (a) Prior to training, the model undergoes iterative inference to reach an equilibrium plan. (b) Then, it is pushed away from the equilibrium towards the ground truth by a supervised loss. This teaches the model to self-refine by mapping a suboptimal equilibrium plan to a better plan.
@@
This new formula represents a new equilibrium sequence modeling scheme that can be implemented in two alternating steps: (1) In the first step, we solve the fixed-point problem by iterative inference of LLM based on the previous output tokens $x_{t}$ until the new output tokens converges, yielding an equilibrium plan $x^{*}$. (2) In the second step, the equilibrium $x^{*}$ is paired with the ground truth plan $y$ as a training sequence, which is used as in the standard supervised finetuning pipeline to teach the LLM to self-refine. The two-step procedure is illustrated in Figure 2.
It features two intuitive advantages: (1) instead of directly regressing the ground truth, it gently adjusts the equilibrium point, which reduces overfitting compared to the vanilla supervised finetuning; (2) by guiding the LLM to map a suboptimal plan $x^{*}$ to a better plan $y$, it teaches self-refinement via a simple supervised loss, without requiring additional value functions or reward models (Welleck et al., 2023; Havrilla et al., 2024; Qu et al., 2024; Kumar et al., 2025).
@@ 3.3 Equilibrium Models with Feedback
To take into account environmental feedback, we consider an adaptive context $c_{t}$ that is influenced by the plan $x_{t}$ rather than fixed. Then, the previously considered equilibrium solving process of Equation 2 should be revised as an iterative process coupling the plan $x_{t}$ with the feedback $c_{t}$, starting from $x_{0}=c_{0}=\varnothing$:
@@
After the modification, the existing derivations only hold when we neglect the derivatives related to $c^{*}$. Fortunately, this is a natural choice due to the non-differentiability of most feedbacks. Therefore, the equilibrium planner can be trained in a similar supervised way as in Equations 6 and 2, and after training it would be able to self-refine based on the latest feedback just by forward passes. However, iteratively interacting with the environment to obtain feedback is costly and may not be recoverable. In response, we devise a nested equilibrium solving scheme for more efficient closed-loop planning.
@@
[…]
Figure 3: Illustration of our proposed framework. It incorporates (1) a memory containing all equilibrium experiences during inference, (2) a self-refining planner trained on equilibrium plans alongside the ground truth, and (3) a world model trained on experiences to simulate environmental feedback. Three modules work synergistically for closed-loop planning.
@@
Thanks to this inner-loop introspection mechanism, our equilibrium model can be more efficient in closed-loop planning, outperforming with fewer environmental interactions.
@@
which corresponds to the starting point $x_{t+1}^{0}$ of the inner loop. Similarly, history feedback could be reused across different inner loops. This is achieved by initializing the context of the next inner loop by concatenating the previous feedback and the latest feedback (paired with its plan):
@@
where $\parallel$ denotes concatenation. The nested inference procedure with reuse of equilibrium is described in Algorithm 1.